OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 1 b) x = 2 c) x = 7 d) x = 5

OpenStudy (amistre64):

x2 + 4x – 5 = 7 x2 + 4x – 12 = 0 x = 6,+2; id pick b: x=2

OpenStudy (amistre64):

dbl chk: 4+8-5=7 12-5=7 7=7

OpenStudy (anonymous):

SO its c'?

OpenStudy (amistre64):

no.... x = 2; that was just me double checking to make sure

OpenStudy (anonymous):

Oh okaye , sorry . thank you

OpenStudy (amistre64):

Gnite :)